# Accelerating Vision Diffusion Transformers with Skip Branches

### v1

@@ Abstract
Diffusion Transformers (DiT), an emerging image and video generation model architecture, has demonstrated great potential because of its high generation quality and scalability properties. Despite the impressive performance, its practical deployment is constrained by computational complexity and redundancy in the sequential denoising process. While feature caching across timesteps has proven effective in accelerating diffusion models, its application to DiT is limited by fundamental architectural differences from U-Net-based approaches. Through empirical analysis of DiT feature dynamics, we identify that significant feature variation between DiT blocks presents a key challenge for feature reusability. To address this, we convert standard DiT into Skip-DiT  with skip branches to enhance feature smoothness. Further, we introduce Skip-Cache  which utilizes the skip branches to cache DiT features across timesteps at the inference time. We validated effectiveness of our proposal on different DiT backbones for video and image generation, showcasing skip branches to help preserve generation quality and achieve higher speedup. Experimental results indicate that Skip-DiT achieves a 1.5×1.5$\times$1.5 × speedup almost for free and a $\times2$ speedup with only a minor reduction in quantitative metrics. Code is available at https://github.com/OpenSparseLLMs/Skip-DiT.git.
@@ 1 Introduction
Numerous approaches have been proposed to enhance the efficiency of diffusion models, including reduced sampling techniques [33], distillation methods [41, 29], and quantization strategies [5]. Caching mechanisms, which reuse noise latents across timesteps, have emerged as a particularly promising direction as they do not require extensive model retraining [18, 15, 38, 47, 6]. However, many existing caching approaches [18, 38] are specifically tailored to U-Net architectures, leveraging their unique structural properties—particularly the skip connections between paired downsampling and upsampling blocks that enable high-level feature reuse while selectively updating low-level features. While some recent studies [47, 6] have attempted to adapt caching mechanisms for DiT acceleration, they have not achieved the same level of efficiency gains and performance preservation as their U-Net counterparts.
@@
Figure 2: Feature smoothness analysis of DiT with class-to-video generation on Taichi using DDPM. Normalized disturbances (with strength coefficients $\alpha$ and $\beta$) are added to the model with and without skip connections. We compare the similarity between the original and perturbed features. The feature difference surface of Latte, with and without skip connections, is visualized at step 10 and 250 of DDPM.
@@
  • We build Skip-DiT , a skip-branch augmented DiT architecture that enhances feature smoothness, and Skip-Cache that efficiently leveraging skip branches for feature caching across timesteps. Skip-Cache facilitates accelerated inference via caching while maintaining visual generation performance.
@@ 2 Related Works
Diffusion model has become the dominating architecture for image and video generation, whose main idea is iterative generate high-fidelity images or video frames from noise [26]. Early diffusion models mainly employ U-Net as their denoising backbone [24, 3]. However, U-Net architectures struggle to model long-range dependencies due to the local nature of convolutions. Researchers proposing diffusion transformer model (DiT) for image generation [4, 2, 23]. Recent years have witnessed a significant growth in studies of video DiT. Proprietary DiT such as Sora [20] and Movie-Gen [25] show impressive generation quality, also evidenced by open-sourced implementation [48, 12]. Latte decomposes the spatial and temporal dimensions into four efficient variants for handling video tokens, allowing effective modeling of the substantial number of tokens extracted from videos in the latent space [19]. CogvideoX adds a 3D VAE combined with an expert transformer using adaptive LayerNorm, which enables the generation of longer, high-resolution videos [40]. However, as the number of tokens grows exponentially with video length and spatial resolution, the computational complexity of DiT especially the self-attention mechanism remains a significant bottleneck for video generation.
@@
Since Diffusion model involves iterative denoising, caching features across time-steps, model layers, and modules has been found an effective way to save inference computation costs. For U-Net Diffusion, DeepCache [18] and FRDiff [32] exploit temporal redundancy by reusing features across adjacent denoising steps. While other works take a more structured approach by analyzing and caching specific architectural components–Faster Diffusion [15] specifically targets encoder feature reuse while enabling parallel decoder computation, and Block Caching [38] introduces automated caching schedules for different network blocks based on their temporal stability patterns. Recently, cache-based acceleration has also been applied to DiT. PAB [47] introduces a pyramid broadcasting strategy for attention outputs. $\Delta$-DiT [6] proposes adaptive caching of different DiT blocks based on their roles in a generation–rear blocks during early sampling for details and front blocks during later stages for outlines. T-Gate [45] identifies a natural two-stage inference process, enabling the caching and reuse of text-oriented semantic features after the initial semantics-planning stage. While these caching techniques have shown promise, they are primarily limited to inference-time optimization, and there remains significant potential for improving their acceleration factors.
@@ 3.1 Preliminaries
In our study, we consider two types of DiT models processing different visual information: image DiT and video DiT, presented in Figure 3 (a). ❶ Image DiT model follows Vision Transformer, each block contains self-attention, cross-attention, and feed-forward networks (FFN), where the cross-attention module incorporates text and timestep conditions. ❷ Video DiT adopts dual-subblock architecture: the spatial Transformer subblock processes tokens within the same time frame, while temporal subblock manages cross-frame relationships. These blocks alternate in sequence, with cross-attention integrating conditional inputs. A complete video DiT block pairs one spatial and one temporal component in an interleaved pattern, following [19, 47].
@@ 3.3 Skip-DiT: Improving Feature Smoothness
We visualize vanilla DiT feature smoothness in Figure 2 (w/o Skip branches) which is trained and generated on Taichi dataset. We can observe drastic feature change at the two DDPM steps, indicating they are not ideal states to cache features, thus shortening the space for caching. According to the insights from [13] and Diffusion caching study utilizing U-Net residual connection feature [18], we next investigate whether DiT feature smoothness can be improved by minimal modification to its structure to insert residual property, i.e. skip branches.
@@ 4.1 Implementation Details
For the text-to-video task, original Latte is trained on Webvid-10M [1] and Vimeo-2M [36], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid-10M is lower than 512$\times$512 and is predominantly used in the early stages of pre-training [20], we utilize only Vimeo-2M for training Skip-DiT. To align with Latte, we sample 330k text-video pairs from Vimeo-2M. All training data are resized to a resolution of 512$\times$512, with 16 frames per video and a frame rate of 8.
@@ 4.3 Generalize Skip-Cache to image Generation
Table 4 provides a comprehensive comparison of Hunyuan-DiT and various caching methods. Notably, our caching mechanism achieves a 2.28$\times$ speedup without any degradation in FID or CLIP scores. Furthermore, it outperforms all other caching methods in terms of PSNR, LPIPS, and SSIM scores, consistently maintaining the highest performance even with a 1.93$\times$ speedup. These findings underscore the robustness and adaptability of our caching mechanism to image generation tasks.
Table 5: Ablation study on different timestep cached. Total timesteps is 1000. During early timesteps (1000$\rightarrow$700), feature changes rapidly. In the latter timesteps (700$\rightarrow$50) feature dynamics becoming considerably smoother. We only perform Skip-Cache  in caching timesteps We highlight the best metrics in bold.
@@
Table 6: Compitablity of Skip-DiT with other caching methods. All the methods are peformed on Skip-DiT . $n=i$ indicates caching the high-level feature in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.
@@ 4.4 Ablation studies
As shown in Table 6, we extend the existing DiT caching methods to Skip-DiT and observe slight performance improvements. Specially, in $\Delta$-DiT, the middle blocks are cached instead of the front blocks, as discussed in Section 4.1, to leverage Skip-DiT’ U-shaped structure. Taking PAB [47] as an example, it loses 1.15% less in VBench score and achieves noticeably better PSNR and SSIM scores on Skip-DiT compared to Latte, highlighting the potential of Skip-DiT to enhance cache-based methods, and proving the superior caching efficiency of model with better feature smoothness.
@@ 6 Class-to-image Generation Experiments
Peebles and Xie [22] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] 512$\times$512 and 256$\times$256 benchmarks. We add skip branches to its largest model DiT-XL/2 to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution 256$\times$256 from scratch with completely the same expirements setting as DiT-XL/2, and far exceeds DiT-XL/2 with only around 38% of its training cost.
@@ 8 Implementation Details
T-Gates divide the diffusion process into two phases: (1) the Semantics-Planning Phase and (2) the Fidelity-Improving Phase. In the first phase, self-attention is computed and reused every $k$ steps. In the second phase, cross-attention is cached using a caching mechanism. The hyperparameter $m$ determines the boundary between these two phases. For our implementation, we use the same hyperparameters as PAB [47]. Detailed configurations are provided in Table 10.

### v2

@@ Abstract
Diffusion Transformers (DiT), an emerging image and video generation model architecture, has demonstrated great potential because of its high generation quality and scalability. Despite the impressive performance, its practical deployment is constrained by computational complexity and redundancy in the sequential denoising process. While feature caching across timesteps has proven effective in accelerating diffusion models, its application to DiT is limited by fundamental architectural differences from U-Net-based approaches. Through empirical analysis of DiT feature dynamics, we identify that significant feature variation between DiT blocks presents a key challenge for feature reusability. To address this, we convert standard DiT into Skip-DiT  with skip branches to enhance feature smoothness. Further, we introduce Skip-Cache  which utilizes the skip branches to cache DiT features across timesteps at the inference time. We validated the effectiveness of our proposal on different DiT backbones for video and image generation, showcasing skip branches to help preserve generation quality and achieve higher speedup. Experimental results indicate that Skip-DiT achieves a 1.5×1.5$\times$1.5 × speedup almost for free and a $\times2$ speedup with only a minor reduction in quantitative metrics. Code is available at https://github.com/OpenSparseLLMs/Skip-DiT.git.
@@ 1 Introduction
Numerous approaches have been proposed to enhance the efficiency of diffusion models, including reduced sampling techniques [33], distillation methods [41, 29], and quantization strategies [5]. Caching mechanisms, which reuse noise latent across timesteps, have emerged as a particularly promising direction as they do not require extensive model retraining [18, 15, 38, 47, 6]. However, many existing caching approaches [18, 38] are specifically tailored to U-Net architectures, leveraging their unique structural properties—particularly the skip connections between paired downsampling and upsampling blocks that enable high-level feature reuse while selectively updating low-level features. While some recent studies [47, 6] have attempted to adapt caching mechanisms for DiT acceleration, they have not achieved the same level of efficiency gains and performance preservation as their U-Net counterparts.
@@
Figure 2: Feature smoothness analysis of DiT with class-to-video generation on Taichi using DDPM. Normalized disturbances (with strength coefficients $\alpha$ and $\beta$) are added to the model with and without skip connections. We compare the similarity between the original and perturbed features. The feature difference surface of Latte, with and without skip connections, is visualized in steps 10 and 250 of DDPM.
@@
  • We build Skip-DiT , a skip-branch augmented DiT architecture that enhances feature smoothness, and Skip-Cache that efficiently leverages skip branches for feature caching across timesteps. Skip-Cache facilitates accelerated inference via caching while maintaining visual generation performance.
@@ 2 Related Works
The diffusion model has become the dominating architecture for image and video generation, whose main idea is iterative generate high-fidelity images or video frames from noise [26]. Early diffusion models mainly employ U-Net as their denoising backbone [24, 3]. However, U-Net architectures struggle to model long-range dependencies due to the local nature of convolutions. Researchers proposing diffusion transformer model (DiT) for image generation [4, 2, 23]. Recent years have witnessed a significant growth in studies of video DiT. Proprietary DiT such as Sora [20] and Movie-Gen [25] show impressive generation quality, also evidenced by open-sourced implementation [48, 12]. Latte decomposes the spatial and temporal dimensions into four efficient variants for handling video tokens, allowing effective modeling of the substantial number of tokens extracted from videos in the latent space [19]. CogvideoX adds a 3D VAE combined with an expert transformer using adaptive LayerNorm, which enables the generation of longer, high-resolution videos [40]. However, as the number of tokens grows exponentially with video length and spatial resolution, the computational complexity of DiT especially the self-attention mechanism remains a significant bottleneck for video generation.
@@
Since the diffusion model involves iterative denoising, caching features across time-steps, model layers, and modules has been found an effective way to save inference computation costs. For U-Net Diffusion, DeepCache [18] and FRDiff [32] exploit temporal redundancy by reusing features across adjacent denoising steps. While other works take a more structured approach by analyzing and caching specific architectural components–Faster Diffusion [15] specifically targets encoder feature reuse while enabling parallel decoder computation, and Block Caching [38] introduces automated caching schedules for different network blocks based on their temporal stability patterns. Recently, cache-based acceleration has also been applied to DiT. PAB [47] introduces a pyramid broadcasting strategy for attention outputs. $\Delta$-DiT [6] proposes adaptive caching of different DiT blocks based on their roles in a generation–rear blocks during early sampling for details and front blocks during later stages for outlines. T-Gate [45] identifies a natural two-stage inference process, enabling the caching and reuse of text-oriented semantic features after the initial semantics-planning stage. While these caching techniques have shown promise, they are primarily limited to inference-time optimization, and there remains significant potential for improving their acceleration factors.
@@ 3.1 Preliminaries
In our study, we consider two types of DiT models processing different visual information: image DiT and video DiT, presented in Figure 3 (a). ❶ Image DiT model follows Vision Transformer, each block contains self-attention, cross-attention, and feed-forward networks (FFN), where the cross-attention module incorporates text and timestep conditions. ❷ Video DiT adopts dual-subblock architecture: the spatial Transformer subblock processes tokens within the same time frame, while the temporal subblock manages cross-frame relationships. These blocks alternate in sequence, with cross-attention integrating conditional inputs. A complete video DiT block pairs one spatial and one temporal component in an interleaved pattern, following [19, 47].
@@ 3.3 Skip-DiT: Improving Feature Smoothness
We visualize vanilla DiT feature smoothness in Figure 2 (w/o Skip branches) which is trained and generated on Taichi dataset. We can observe drastic feature changes at the two DDPM steps, indicating they are not ideal states to cache features, thus shortening the space for caching. According to the insights from [13] and Diffusion caching study utilizing U-Net residual connection feature [18], we next investigate whether DiT feature smoothness can be improved by minimal modification to its structure to insert residual property, i.e. skip branches.
@@ 4.1 Implementation Details
For the text-to-video task, the original Latte is trained on Webvid-10M [1] and Vimeo-2M [36], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid-10M is lower than 512$\times$512 and is predominantly used in the early stages of pre-training [20], we utilize only Vimeo-2M for training Skip-DiT. To align with Latte, we sample 330k text-video pairs from Vimeo-2M. All training data are resized to a resolution of 512$\times$512, with 16 frames per video and a frame rate of 8.
@@ 4.3 Generalize Skip-Cache to image Generation
Table 4 provides a comprehensive comparison of Hunyuan-DiT and various caching methods. Notably, our caching mechanism achieves a 2.09$\times$ speedup without any degradation in FID or CLIP scores. Furthermore, it outperforms all other caching methods in terms of PSNR, LPIPS, and SSIM scores, consistently maintaining the highest performance even with a 1.96$\times$ speedup. These findings underscore the robustness and adaptability of our caching mechanism to image generation tasks.
Table 5: Ablation study on different timestep cached. Total timesteps is 1000. During early timesteps (1000$\rightarrow$700), the feature changes rapidly. In the latter timesteps (700$\rightarrow$50) feature dynamics become considerably smoother. We only perform Skip-Cache  in caching timesteps We highlight the best metrics in bold.
@@
Table 6: Compatibility of Skip-DiT with other caching methods. All the methods are performed on Skip-DiT . $n=i$ indicates caching the high-level feature in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.
@@ 4.4 Ablation studies
As shown in Table 6, we extend the existing DiT caching methods to Skip-DiT and observe slight performance improvements. Specifically, in $\Delta$-DiT, the middle blocks are cached instead of the front blocks, as discussed in Section 4.1, to leverage Skip-DiT’ U-shaped structure. Taking PAB [47] as an example, it loses 1.15% less in VBench score and achieves noticeably better PSNR and SSIM scores on Skip-DiT compared to Latte, highlighting the potential of Skip-DiT to enhance cache-based methods, and proving the superior caching efficiency of the model with better feature smoothness.
@@ 6 Class-to-image Generation Experiments
Peebles and Xie [22] proposed the first diffusion model based on the transformer architecture, and it outperforms all prior diffusion models on the class conditional ImageNet [7] 512$\times$512 and 256$\times$256 benchmarks. We add skip branches to its largest model DiT-XL/2 to get Skip-DiT. We train Skip-DiT on class conditional ImageNet with resolution 256$\times$256 from scratch with completely the same experiments setting as DiT-XL/2, and far exceeds DiT-XL/2 with only around 38% of its training cost.
@@ 8 Implementation Details
T-Gates divide the diffusion process into two phases: (1) the Semantics-Planning Phase and (2) the Fidelity-Improving Phase. In the first phase, self-attention is computed and reused every $k$ step. In the second phase, cross-attention is cached using a caching mechanism. The hyperparameter $m$ determines the boundary between these two phases. For our implementation, we use the same hyperparameters as PAB [47]. Detailed configurations are provided in Table 10.